pleas provide excel spreadsheet showing calculations thank youRead each question carefully! Make sure that you have answered all parts of each question. Please write your final answers in
Answer #1

Problem 1

Question Answer
Total Revenue Budget $    1,050,000
Production budget 5,100 units
Direct materials budget 20,300 pounds
Direct materials budget $       101,500
Direct labor budget 15,300 hours
Direct labor budget $       336,600
Overhead budget $       336,600
Budgeted GP $       275,300

-Total Revenue Budget = Units sold X Selling price = 5,000 X 210 = $ 1,050,000

-Production budget in units = Target inventory + Sales - Current Inventory = 300 + 5,000 - 200 = 5,100 units

-DIrect material in units = Target inventory + Sales - Current Inventory = 100 + (5,100 X 4) - 200 = 20,300 pounds

-Direct materials in dollars = 20,300 X 5 = $101,500

-Direct labor in hours = 5,100 X 3 = 15,300 hours

-Direct labor in dollars = 15,300 X 22 = 336,600

-Overhead budget in dollars = 5,100 X 3 X 22 = 336,600

-Budgeted Gross Profit = Sales - Material - Labor - Overhead = 1,050,000 - 101,500 - 336,600 - 336,600 = 275,300